OpenStudy (cderaalien):

Don't know how to solve, tried elimination and substitution!! v + c = 27 4v + c = 45

OpenStudy (anonymous):

wait what are u solving for?

OpenStudy (anonymous):

v or c?

OpenStudy (alexandervonhumboldt2):

this is so easy

OpenStudy (alexandervonhumboldt2):

c=27-v c=45-4v

OpenStudy (anonymous):

yes for u @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

equal: 27-v=45-4v solve for v

OpenStudy (anonymous):

well done ur soo soos ooosoo smart !

OpenStudy (cderaalien):

......

OpenStudy (willie579):

@AlexandervonHumboldt2 Liek a baws. Gud job.

OpenStudy (anonymous):

atleast let him try first himself

OpenStudy (cderaalien):

*her

OpenStudy (cderaalien):

But okay sure let me try

OpenStudy (cderaalien):

With the problems above, can I use elimination or substitution, or both?

OpenStudy (anonymous):

yes u can use both

OpenStudy (alexandervonhumboldt2):

yeah both are useful

OpenStudy (cderaalien):

Ah! v = 24

OpenStudy (cderaalien):

Now, I plug into one of the original equations, correct?

OpenStudy (anonymous):

yes

OpenStudy (alexandervonhumboldt2):

waittttttttttttttttttttttt

OpenStudy (alexandervonhumboldt2):

27-v=45-4v 3v=18 v=6.....................

OpenStudy (cderaalien):

lmao...oh god

OpenStudy (willie579):

O_O http://prntscr.com/9goc4g

OpenStudy (alexandervonhumboldt2):

aweeeeeee

OpenStudy (alexandervonhumboldt2):

wait ima so stupid

OpenStudy (cderaalien):

I started by adding 27 to both sides.

OpenStudy (anonymous):

lol i thought u did right srry i didnt check cause @AlexandervonHumboldt2 already did it on top

OpenStudy (alexandervonhumboldt2):

sorry let me start as i did:

OpenStudy (alexandervonhumboldt2):

v + c = 27 4v + c = 45 => c=27-v c=45-4v => 45-4v=27-v v=6 thus: c=27-6=21 thus 6 and 21

OpenStudy (alexandervonhumboldt2):

yeah sorry you was correct

OpenStudy (willie579):

*were

OpenStudy (alexandervonhumboldt2):

lol willie

OpenStudy (cderaalien):

Oh, so 24 = v is correct?

OpenStudy (anonymous):

yes it is , i got 24 too

OpenStudy (anonymous):

u just plug into any eq v=24

OpenStudy (cderaalien):

Okay so for example, v + c = 27 I would do it like this, 24 + c = 27 Subtract 24 from both sides, which means c = 3?

OpenStudy (cderaalien):

Do we all agree?

OpenStudy (anonymous):

yes well done

OpenStudy (cderaalien):

Thank you! All of you!!

OpenStudy (shadowlegendx):

@AlexandervonHumboldt2 This is so easy http://prntscr.com/9gofbi ._.

OpenStudy (alexandervonhumboldt2):

LOL

OpenStudy (alexandervonhumboldt2):

YOU ARE A COOKIE EVIL! YOU DIDN'T BET RESPONDED ME XD HAHAA JOKE

OpenStudy (alexandervonhumboldt2):

@CookieEvilX

OpenStudy (alexandervonhumboldt2):

lol